The second of three numbers is 8 more that the first and the third number is 3 less than 3 times the first. If the tird number is 15 more than the second, find the three numbers.
Here's what I have for the variables:
First number = x
Second number = x + 8
third number = 3x - 3 and it is x + 8 + 15
Where do I go from here??

Answer by venugopalramana(3286)

The second of three numbers is 8 more that the first and the third number is 3 less than 3 times the first. If the tird number is 15 more than the second, find the three numbers.
Here's what I have for the variables:
First number = x
Second number = x + 8
third number = 3x - 3 and it is x + 8 + 15
Where do I go from here??
OH!! DEAR!! DEAR!! YOU GOT IT !! YOU ARE ALMOST AT THE FINISH LINE !!! JUMP AHEAD ! YOU WILL GET THE ANSWER!!!
You rightly concluded that the third number is 3x-3 and it is x+8+15...MEANS WHAT ???? THEY ARE EQUAL ...SO
3x-3=x+8+15
3x-x=8+15+3
2x=26...or x=13..... hence the 3 numbers are as per your own working
First number = x =13
Second number = x + 8=13+8 =21
third number = 3x - 3=3*13-3=36..check this if you want with
x + 8 + 15 = 13+8+15 =36